each of the following: to 2 b. c.
Answer #1

(a) Observed time is the time measured to complete the task. So it is 1.21 minutes.

(b) Normal Time = Average Observed Time x Performance Rating Factor

Normal Time = 1.21 x 91% = 1.10 minutes

(c)

\textup{Standard Time =}\frac{Total Normal Time}{1-Allowance factor}

\textup{Standard Time =}\frac{1.10}{1-0.05}=1.16 \mathbf{\textup{ minutes}}